OpenStudy (anonymous):

Estimate the product of 63,478 and 235

OpenStudy (sugarmochi):

you estimate it give a guess on how much those two numbers equal

OpenStudy (sugarmochi):

there is no write and wrong answer for estimating

ganeshie8 (ganeshie8):

if you love zeroes : \[\large 60,000 \times 200\]

OpenStudy (anonymous):

yup.... but m not sure how to do it...

OpenStudy (anonymous):

but ok thx ganeshie8

OpenStudy (anonymous):

thx @ganeshie8 and @SugarMochi

ganeshie8 (ganeshie8):

simply multiply 6*2 = 12 and append the 0's in the end : 12,000,000

ganeshie8 (ganeshie8):

yw!

OpenStudy (mathmate):

For those who are used to accurate estimating in the head, we can take 250 instead of 235: 64000*250=64000*1000/4=16000000 However, 235 equals 250 less 6%, so subtract 6%=1000000 to get 15000000 as a better estimate.